Suppose that a firm’s production function is given by Q = KL + K, with MPK = L + 1 and MPL = K. At point A, the firm uses K = 3 units of capital and L = 5 units of labor. At point B, along the same isoquant, the firm would only use 1 unit of capital. a) Calculate how much labor is required at point B.
1
Expert's answer
2019-10-01T12:33:47-0400

a) At point A: Q = 3×5 + 3 = 18 units.

If at point B, along the same isoquant, the firm would only use 1 unit of capital, then Q = 1L + 1 = 18,

L = 17 units of labor is required at point B.